# Design of a Quality Management System based on the EU Artificial Intelligence Act

### v1

@@ 3.1 Requirements
The QMS comprises functional (FR) and non-functional (NFR) requirements grouped into three distinct types. As defined in [6], FRs describe the system’s functionalities, such as the feature of evaluating an AI system’s performance. In contrast, NFRs describe how the system should perform the functionality, such as ensuring that the AI system’s performance evaluation is done in under 2 seconds. Firstly, the QMS incorporates legal requirements derived from interpreting legal regulations set out in the EU AIA for high-risk AI and GPAI systems to ensure comprehensive safety and compliance. Secondly, system design requirements specify the technical aspects of the QMS. These requirements cover human involvement, architecture, and computational needs. All requirements are high-level and will be subdivided into more concrete sub-requirements in future iterations. However, they represent the minimum necessary functionality for such a QMS and illustrate the extensive range of features that need to be integrated.
Legal Requirements. Several requirements are derived from the given legal regulations. Articles 53 and 55 EU AIA provide information on requirements specifically for GPAI models, shown in Table LABEL:tab:LR01.
@@
The GPAI-specific requirements align closely with the requirements constructed based on the regulations for high-risk AI systems. The idea is to integrate all requirements for high-risk AI systems into the QMS. According to the EU AIA, the QMS  [^11] Check for details and all obligations: Chapter 3, Section 3 Article 17 Quality Management System EU AIA should generally encompass strategies for regulatory compliance, including conformity assessment and management for modifications of legal guidelines or technical features. It should also provide techniques or procedures for the model’s design, design control, verification, quality control, and quality assurance. Specifically, the QMS should contain an RMS to identify, analyze, assess, and mitigate potential risks and a DMDGS to demonstrate the data quality used for training, validation, and testing. The results should be stored in technical documentation. The software requirements derived from the obligation of high-risk AI systems are listed in Table LABEL:tab:LR02 and must be ensured within the QMS.
@@
System Design Requirements. The system design requirements are categorized into three distinct types. First, The User Interface (UI) design and interaction specifications should encourage human involvement in checking and documenting AI systems. Second, architectural requirements should ensure a modular design, a smooth flow of data and communication between the services, and easy maintenance. The design should apply to multiple AI systems across various domains and tasks. Third, computational requirements should guarantee the efficient execution of technical evaluation metrics, even for large AI systems such as LLMs, to maintain optimal performance. According to [28], the risk assessment process should adopt a human-centered design approach. Moreover, besides human involvement in the RMS, users should be able to include references and descriptions of the data utilized for training, validation, and testing, consolidating all necessary information required by legal standards within the DMDGS. Additionally, users should be able to directly access and download technical documentation from the QMS to verify the test and validation proceedings on the AI system and ensure compliance with EU AIA regulations. Integrating all these functionalities into several sub-modules within a single tool aims to reduce effort, time, and costs for AI system providers and deployers. Table LABEL:tab:SDR01 lists all UI and human involvement requirements.
@@
The QMS should have a modular design to react quickly to future changes in the legislative framework. The modular design is ensured by designing independent sub-services for each article in the EU AIA, which can be plugged into the QMS and maintained and updated separately at any time. Key modules include the RMS (cf. Article 9 EU AIA) and the DMDGS (cf. Article 10 EU AIA), which are integrated into the first version of the prototype QMS. Potential additional modules, such as the AI system’s event logging (cf. Article 12 EU AIA), can be added in future work. The modules should be designed to apply generically to any type and architecture of AI system, except for the specific technical evaluation metrics, which need to align with the type and architecture of the underlying AI system. Additionally, the QMS allows users to persistently store technical documentation, past risk assessment processes, and data check references. This feature not only ensures comprehensive record-keeping but also facilitates easy access to historical data, which is crucial for maintaining compliance and continuous improvement. The goal is to enable the QMS to be utilized throughout the entire lifecycle of the AI system, including post-market. All requirements of the system design architecture are detailed in Table LABEL:tab:SDR02.
@@
Identifying the required computational resources is the third type of system design requirement. LLMs, for example, require significant computational power for tasks like calculating the gradients needed for some technical evaluation metrics. These gradients are necessary for certain adversarial attacking and explainability techniques. The specific GPU and CPU resources need to be tested, evaluated, and defined for a second version of the prototype QMS. Table LABEL:tab:SDR03 details the computational requirements for the system design.
@@ 3.2 Architecture and Data Models
The QMS architecture is based on a microservice design (cf. [25]). The current prototype QMS consists of an RMS module, a dedicated DMDGS module, and a user authentication module, which only stores user and login information. All modules are independent services containing their backend and database. The architecture of the QMS and the data models for the RMS, DMDGS, and user authentication databases as UML class diagrams are illustrated in Fig. 1.
@@
Users can interact with the UI implemented on the frontend to participate in the verification and documentation process within the QMS by executing risk assessments and ensuring data management and governance compliance. An API Gateway backend service orchestrates all user requests from the frontend and forwards each request to the corresponding backend system. The API Gateway loads an environment file containing the names or pseudonyms for each sub-service root. The backend of the corresponding sub-service then sends the response back to the API Gateway, which forwards the message to the frontend, where it is loaded into the UI for the user. This structure, data transfer, and orchestration reduce complexity and improve modularity. Complexity is reduced, and modularity is ensured because adding sub-services to the QMS does not alter existing data transfers. To add a new service to the communication and data transfer pipeline, only an additional entry in the environment file read by the API Gateway is required. This prototype QMS focuses on the functionalities of the RMS rather than those in the DMDGS, as already mentioned. The RMS backend consists of two different components. It contains a verification component that can load different language models available from Hugging Face and implements all technical evaluation metrics presented in the next section (cf. Sect. 3.3), which are used to analyze and quantitatively assess the LLM’s performance, consistency, and explainability for specialized tasks in selected high-risk domains. Furthermore, the RMS backend contains a communication component responsible for frontend-backend-database communication and data transfer. Each sub-service, including the DMDGS and user authentication service, contains a communication component with the same structure and functionality. In contrast to the RMS, the DMDGS, and the user authentication services only include a communication component and no further components. The RMS data model consists of three main classes: risk identification, analysis, and assessment. A risk assessment object stores one risk identification and one risk analysis object. Users can execute multiple risk assessments, with the risk assessment IDs stored in the user object. This setup ensures that users can quickly review past risk assessment results in the UI. The DMDGS data model follows the same principle. The user class stores a list of data IDs, where each data object is linked to an LLM object and a data check object that verifies the compliance of the added data reference object. It is important to note that the RMS risk analysis requires data, which has a domain and task to assess the LLM’s characteristics, efficiency, and risks (performance, explainability, and consistency). On the other hand, the data referenced in the DMDGS are the training, validation, and testing data to train and develop the LLM which will be assessed in the RMS after it is ready to use. The user authentication data model only contains a user class to store user IDs, personal data, and login information. This database encrypts and hashes sensitive data and is made secure as the only database storing sensitive personal data. Whenever a new user signs up, the user is stored in the database, and the created user ID (random string created by MongoDB) is distributed to each user collection in the other sub-services databases, which ensures consistency, with each user collection storing identical user IDs. The following technology stack is used to implement the prototype QMS: The frontend is developed using JavaScript and the React.js  [^12] React.js: https://react.dev, accessed on 7 May 2024. library. All backend services are written in Python, utilizing various libraries and packages such as PyTorch  [^13] PyTorch: https://pytorch.org, accessed on 7 May 2024., Transformers  [^14] Transf.: https://huggingface.co/docs/transformers/, accessed on 15 May 2024. in the verification component to i) load the LLMs from Huggingface into the GPU and to ii) perform computations on the loaded LLMs, and FastAPI  [^15] FastAPI: https://fastapi.tiangolo.com, accessed on 7 May 2024. and PyMongo  [^16] PyMongo: https://pymongo.readthedocs.io/en/, accessed on 24 July 2024. to implement all REST and CRUD methods to provide a communication pipeline from the frontend to the backend and from the backend to the database. MongoDB  [^17] MongoDB: https://www.mongodb.com, accessed on 7 May 2024., a non-SQL database, is employed for data storage, allowing for rapid and straightforward modifications and design changes. MongoDB uses so-called collections (database tables) which can be modeled and designed using UML class diagrams and can always be extended or changed, keeping them flexible. The documents are the elements of a collection (rows of a table) and are stored in the collections as JSON data, maintaining a unified structure and language used in the frontend. A document can contain different collections of document data, similar to object-oriented programming, making MongoDB easy to understand.
3.3 Prototype QMS - Version 1
The first version of the prototype QMS can be accessed under the following URL: https://power.bpm.cit.tum.de/qmsAIA/.
Main Service: Quality Management System. The user accesses the QMS home page after signing up and signing in. The prototype QMS home page design is depicted in Fig. 2. To sign in, the user must enter his/ her username, email address, and password. If the user has no account, he/ she can click on the sign-up button to access the sign-up page and create a new account. The homepage is subdivided and structured into the prototype’s several sub-services. Each functional sub-service (RMS and DMDGS) is listed in a separate row, one below the other, and contains two boxes each, one for performing risk assessments or data checks and one for viewing past risk assessments or data checks. The user authentication sub-service contains no UI components and is not classified as a functional sub-service, as its aim is solely to securely store user data. Additionally, the UI provides access to various sections with details about the LLMs, a page to add verification data for risk assessments, and relevant about the EU AIA regulations. All these pages can be accessed through the buttons on the tab bar in the header.

### v2

@@ 3.1 Requirements
The QMS comprises functional (FR) and non-functional (NFR) requirements grouped into three distinct types. As defined in [6] (pp. 101-102), FRs describe the system’s functionalities, such as evaluating the AI system’s performance. In contrast, NFRs describe how the system should perform the functionality, such as ensuring that the AI system’s performance evaluation is done in under 2 seconds. The QMS incorporates Legal Requirements (LR) derived from interpreting legal regulations set out in the AIA for high-risk AI and GPAI systems. Furthermore, System Design Requirements (SDR) specify technical aspects of the QMS. These requirements cover human involvement, architecture, and computational and memory needs. All requirements are high-level and will be subdivided into more concrete sub-requirements in future iterations. However, they represent the minimum necessary functionality for such a QMS and illustrate the extensive range of features that need to be integrated.
Legal Requirements. Several requirements are derived from the given legal regulations. Articles 53 and 55 AIA provide information on requirements specifically for GPAI models, shown in Table LABEL:tab:LR01.
@@
The GPAI-specific requirements align closely with the requirements constructed based on the regulations for high-risk AI systems. The idea is to integrate all requirements for high-risk AI systems into the QMS. According to the EU AIA, the QMS should generally encompass strategies for regulatory compliance, including conformity assessment and management for modifications of legal guidelines or technical features. It should also provide techniques or procedures for the model’s design, design control, verification, quality control, and quality assurance. Specifically, the QMS should contain an RMS to identify, analyze, assess, and mitigate potential risks and a DMDGS to demonstrate the data quality used for training, validation, and testing. The results should be stored in technical documentation. The software requirements derived from the obligation of high-risk AI systems are listed in Table LABEL:tab:LR02 and must be ensured within the QMS.
@@
System Design Requirements. The SDRs are categorized into three distinct types. First, The User Interface (UI) design and interaction specifications should encourage human involvement in checking and documenting AI systems. Second, architectural requirements should ensure a modular design, a smooth flow of data and communication between the services, and easy maintenance. The design should apply to multiple AI systems across various domains and tasks. Third, computational requirements should guarantee the efficient execution of technical evaluation metrics, even for large AI systems such as LLMs, to maintain optimal performance. According to [37], the risk assessment process should adopt a human-centered design approach. Moreover, besides human involvement in the RMS, users should be able to include references and descriptions of the data utilized for training, validation, and testing, consolidating all necessary information required by legal standards within the DMDGS. Additionally, users should be able to directly access and download technical documentation from the QMS to verify the test and validation proceedings on the AI system and ensure compliance with AIA regulations. Integrating all these functionalities into several sub-modules within a single tool aims to reduce effort, time, and costs for AI system providers and deployers. Table LABEL:tab:SDR01 lists all UI and human involvement requirements.
@@
The QMS should have a modular design to react quickly to future changes in the legislative framework. The modular design is ensured by designing independent sub-services for each article in the AIA, which can be plugged into the QMS and maintained and updated separately at any time. Key modules include the RMS (cf. Art. 9 AIA) and the DMDGS (cf. Art. 10 AIA), which are integrated into the first version of the prototype QMS. Potential additional modules, such as the AI system’s event logging (cf. Art. 12 AIA), can be added in future work. The modules should be designed to apply generically to any type and architecture of AI system, except for the specific technical evaluation metrics, which need to align with the type and architecture of the underlying AI system. Additionally, the QMS allows users to persistently store technical documentation, past risk assessment processes, and data check references. This feature ensures comprehensive record-keeping and facilitates easy access to historical data, which is crucial for maintaining compliance and continuous improvement. The goal is to enable the QMS to be utilized throughout the entire lifecycle of the AI system, including post-market. All requirements of the system design architecture are detailed in Table LABEL:tab:SDR02.
@@
Identifying the required computational resources is the third type of system design requirement. LLMs, for example, require significant computational power for tasks like calculating the gradients needed for some technical evaluation metrics. These gradients are necessary for certain adversarial attacking and explainability techniques. The specific GPU and CPU resources must be tested, evaluated, and defined for a second version of the prototype QMS. Table LABEL:tab:SDR03 details the computational requirements for the system design.
@@
3.2 Technical Evaluation Metrics
Conducting technical metrics on the AI system is essential to comply with the AIA regulations, such as shown by [13] for Art. 9 AIA to assess risks, or [17] for Art. 14 AIA testing the AI system’s transparency and mandatory in Art. 15 AIA to assess the AI system’s accuracy and robustness. Since the first version of the prototype connects to an LLM, technical metrics for neural networks, specifically LLMs, are defined and implemented. Several metrics already exist to assess classification or regression models on performance, explainability, robustness, or fairness, as presented in works by [13, 34, 4]. However, evaluating generative models on these characteristics presents unique challenges. The technical metrics presented are specifically designed to evaluate the performance, explainability, and robustness of the connected Phi-3-mini-128k-instruct on specialized tasks. The proposed specialized metrics are optimized for text summarization and transformation tasks. This selection aligns with the application of LLMs in automated compliance verification approaches in previous works (cf. [3, 30, 31]), which could have a strong use case in potential high-risk domains such as the legal and financial domains. This selection has been chosen as an example based on previous research. Any other tasks and technical metrics could have been chosen.
Basic Definitions. A text has always a domain and a type, for example, a legal text in the medical domain. The text $t$ consists of $n\in\mathbb{N}$ tokens: $t=t_{0},\ldots,t_{n-1}$. The LLM calculates with numerical values. For that, each token is encoded and embedded in a numerical vector which is the input for LLM inferences. Hence, the encoding function and decoding functions are defined as:
$enc\colon$ $\text{token}\to\mathbb{R}^{\text{embed dim}},$ $dec\colon$ $\mathbb{R}^{\text{embed dim}}\to\text{token}.$
The encoded tokens are then defined as: $enc(t)=x$. An LLM is defined as a function $f$ that takes a sequence of $n$ embedded tokens and returns a sequence of $m$ embedded tokens:
$f\colon\mathbb{R}^{n\times\text{embed dim}}\to\mathbb{R}^{m\times\text{embed dim}},\;x\mapsto\tilde{y}$
The logit values $\tilde{y}$ outputted by the LLM can be decoded to obtain the generated output tokens: $dec(\tilde{y})=\hat{y}$. Additionally, each input text has a predefined ground-truth target output: ${y}$.
Performance. According to [2], one generic and always efficient performance metric is calculating the LLM’s accuracy score for a specialized task. Since we concentrate on the evaluation of summarization and text transformation, true positives (TP) are defined as a set of actor-activity pairs that are present in the predicted output text $\hat{y}$ and the ground truth text $y$. True negatives are defined as a set of actor-activity pairs that are not present in $\hat{y}$ and not in $y$. False positives are defined as a set of actor-activity pairs that are not present in $\hat{y}$ but in $y$. False negatives are defined as a set of actor-activity pairs that are present in $\hat{y}$ but not in $y$. Furthermore, as noted in [2], the BLEU score (described in [38]) is typically employed for assessing language model performance in translation tasks. In contrast, as detailed in [29], the ROUGE score is the more suitable metric for language model performance assessment in summarization tasks. Several ROUGE methods exist. The chosen one here is the ROUGE-N method.
$\operatorname{ROUGE-N}=\frac{|\operatorname{n-gram}(y)\cap\operatorname{n-gram}(\hat{y})|}{|\operatorname{n-gram}(y)|}$
A n-gram is defined as: $\operatorname{n-gram}(t_{0},...,t_{m-1})=\{t_{i},t_{i+1},...,t_{i+n-1}\mid 0\leq i\leq m-n\}$. Here, we defined the recall ROUGE-N, but the precision pendant and the F1 ROUGE-N score are also implemented. The following gives an example of how the n-grams are constructed.
$t=$ The user creates a new process instance, then the system can execute the instance and stop it afterward. $\hat{y}=$ [“User”, “creates new process instance”, “system”, “execute instance”, “system”, “stops instance”]. $\operatorname{2-gram}(\hat{y})=$ [“User creates new process instance”, “creates new process instance system”, “system execute instance”, “execute instance system”, “system stops instance”]. Another implemented performance metric, explained in [20], which is a general performance metric for neural network-based architectures, is the Perplexity score. It computes the uncertainty of the LLM for predicting each subsequent token in the output sequence. A lower perplexity value signifies less uncertainty and, thus, a higher confidence in the correctness of the prediction.
$\text{PP}(f,t)=\left(\prod_{i=0}^{n-1}\frac{1}{P(\tilde{y}_{i}|\tilde{y}_{i-1},\dots,\tilde{y}_{0})}\right)^{\frac{1}{n}}$
Thereby, the $P(\tilde{y}_{i}|\tilde{y}_{i-1},\dots,\tilde{y}_{0})$, is the conditional probability of each next token outputted by the LLM.
Explainability. The explainability of black-box models such as neural networks and, more specifically, huge neural networks such as LLM is still a hot topic in ongoing research. However, three distinct general explainability approaches exist for LLM. Perturbation-based, gradient-based, and self-explanation-based methods. Examples of perturbation-based methods are occlusion and LIME [21, 27]. Occlusion measures the difference between a generated LLM output based on the entire input sequence and the output based on the same input sequence with one word deleted [28]. LIME removes a random subset of words from the input sequence, then computes the model’s output and compares the output with the model’s outputs by removing just one word from the output. They compute a linear regression model comparing the relationship between the different input and output types [40]. Secondly, gradient-based explainability methods exist. One method originally invented for image classifiers by [43] is the generation of saliency maps. Originally, saliency maps aimed to visualize and rank pixels of an input image based on their sensitivity and influence on the predicted output class. [27] also employed the gradient-based approach. They devised a post-hoc explanation technique by formulating a meta-prompt using the most significant input tokens to produce the specified output. This method aims to enhance prompting language models by comprehending the model’s reasoning behind generating a particular output. It involves computing the gradients of the LLM’s target output and each input token and selecting the top-k tokens that exerted the most influence on the model output. The third approach involves LLM self-explanation methods. [21] delineate two distinct types of LLM self-explanation. Firstly, the explanation-prediction method requires the model to identify the most pertinent tokens in the input sequence and subsequently generate a prediction based on them. Secondly, the prediction-explanation method necessitates the model to provide a prediction and then elucidate the rationale behind that prediction. This becomes interesting for models like GPT-4, which are not openly available, making it impractical to compute perturbations and gradients.
In this prototype, a gradient-based saliency map metric is implemented since gradient-based methods are the most promising in getting the hard facts on how the model processes the input.
$S={\frac{\delta\tilde{y}}{\delta x}}=\{\sigma_{0},\dots,\sigma_{n-1}\}$
Here, $S$ denotes the set of gradient values for each input token based on the model output. In a second step, all computed gradient values are normalized: $\lVert S\rVert=\{\lVert\sigma_{0}\rVert,\dots,\lVert\sigma_{n-1}\rVert\}$. Each normalized gradient value ranges from 0 to 1, with values closer to 1 exerting a stronger sensitivity and therefore influence on the predicted output than those at 0. Utilizing these normalized gradient values, the corresponding input tokens are visualized. Tokens with greater influence are depicted in lighter shades of green, while those with lesser influence appear in darker shades of red. This approach produces a token-based saliency map, providing users insight into which input tokens have the most significant impact on the generated output.
Consistency. In this prototype, a gradient-based adversarial examples metric is chosen to be implemented as a consistency method. The overall idea is to measure the degree and effort of changes in the input text required to generate a predetermined false output. This concept is related to generating gradient-based adversarial examples (cf. [46, 16]) and is a robustness and security check for neural networks. The consistency metric is based on a gradient-based adversarial example creation for LLM, similar to the concept introduced by [51]. This approach utilizes adversarial example creation to induce hallucinations in the LLM’s output, effectively evaluating its robustness. Mathematically, it is similar to the gradient-based explainability metric. It involves modifying the text input data by modifying the top-k most important tokens until the LLM generates with high certainty a similar or even the predetermined false output containing hallucinations. The robustness, reliability, and security of the LLM can be evaluated by analyzing the type and quantity of replaced words and the number of replacement iterations. First, the gradient values $S$ are calculated for each token as in the presented gradient-based explainability metric. In the second step, the input tokens are modified by a factor $\eta$, which can be set and adjusted, similar to the learning rate in neural network optimization. The factor must be added in the direction of the gradient value. The goal is to modify the input in a direction to fool the LLM such that the lowest loss for $f(x)$
$x=\sum_{i=0}^{n-1}x_{i}+\eta\cdot\operatorname{sign}(\sigma_{i})$
The embedded input token modification is done once the LLM outputs the predetermined adversarial target. The model’s input from the previous iteration is modified in each iteration to calculate the gradients.
Gradient-based metrics were chosen for explainability and consistency (robustness) due to several important factors. These metrics are fundamental for understanding a model’s output and sensitivity to input data. They offer crucial insights into the model’s architecture and reasoning. Additionally, gradient-based metrics are among the most challenging to implement and compute, especially considering the required GPU memory. This aspect was particularly relevant for the prototype, as it allowed for the analysis of the system’s computational and memory needs.
@@ 3.3 Architecture and Data Models
The QMS architecture is based on a microservice design (cf. [33]). The current prototype QMS consists of an RMS module, a dedicated DMDGS module, and a user authentication module, which only stores user and login information. All modules are independent sub-services containing their backend and database. The architecture of the QMS and the data models for the RMS, DMDGS, and user authentication databases as UML class diagrams are illustrated in Fig. 2.
@@
Users can interact with the UI implemented on the frontend to participate in the verification and documentation process within the QMS by executing risk assessments and ensuring data management and governance compliance. An API Gateway backend service orchestrates all user requests from the frontend and forwards each request to the corresponding backend system. The API Gateway loads an environment file containing the names or pseudonyms for each sub-service root. The backend of the corresponding sub-service then sends the response back to the API Gateway, which forwards the message to the frontend, where it is loaded into the UI for the user. This structure, data transfer, and orchestration reduce complexity and improve modularity. Complexity is reduced, and modularity is ensured because adding sub-services to the QMS does not alter existing data transfers. To add a new service to the communication and data transfer pipeline, only an additional entry in the environment file read by the API Gateway is required. This prototype QMS focuses on the functionalities of the RMS rather than those in the DMDGS, as already mentioned. The RMS backend consists of two different components. It contains a verification component that can load different language models available from Hugging Face and implements all technical evaluation metrics presented in the next section (cf. Sect. 3.4), which are used to analyze and quantitatively assess the LLM’s performance, consistency, and explainability for specialized tasks in selected high-risk domains. Furthermore, the RMS backend contains a communication component responsible for frontend-backend-database communication and data transfer. Each sub-service, including the DMDGS and user authentication service, contains a communication component with the same structure and functionality. In contrast to the RMS, the DMDGS, and the user authentication services only include a communication component and no further components. The RMS data model consists of three main classes: risk identification, analysis, and assessment. A risk assessment object stores one risk identification and one risk analysis object. Users can execute multiple risk assessments, with the risk assessment IDs stored in the user object. This setup ensures that users can quickly review past risk assessment results in the UI. The DMDGS data model follows the same principle. The user class stores a list of data IDs, where each data object is linked to an LLM object and a data check object that verifies the compliance of the added data reference object. It is important to note that the RMS risk analysis requires data, which has a domain and task to assess the LLM’s characteristics, efficiency, and risks (performance, explainability, and consistency). On the other hand, the data referenced in the DMDGS are the training, validation, and testing data to train and develop the LLM, which will be assessed in the RMS after it is ready to use. The user authentication data model only contains a user class to store user IDs, personal data, and login information. The following technology stack is used to implement the prototype QMS: The frontend is developed using JavaScript and the React.js  [^11] React.js: https://react.dev, accessed on 17 October 2024. library. All backend services are written in Python, utilizing various libraries and packages such as PyTorch  [^12] PyTorch: https://pytorch.org, accessed on 17 October 2024, Transformers  [^13] Transf.: https://huggingface.co/docs/transformers/, accessed on 17 October 2024 in the verification component to load the LLMs from Huggingface into the GPU and to perform computations on the loaded LLMs, and FastAPI  [^14] FastAPI: https://fastapi.tiangolo.com, accessed on 17 October 2024 and PyMongo  [^15] PyMongo: https://pymongo.readthedocs.io/en/, accessed on 17 October 2024 to implement all REST and CRUD methods to provide a communication pipeline from the frontend to the backend and from the backend to the database. MongoDB  [^16] MongoDB: https://www.mongodb.com, accessed on 17 October 2024, a non-SQL database, is employed for data storage, allowing for rapid and straightforward modifications and design changes. MongoDB uses so-called collections (database tables), which can be modeled and designed using UML class diagrams and can always be extended or changed, keeping them flexible. The documents are the elements of a collection (rows of a table). They are stored in the collections as JSON data, maintaining a unified structure and language used in the frontend. A document can contain different collections of document data, similar to object-oriented programming, making MongoDB easy to understand.
RMS Verification Component. A special design approach is chosen to implement the technical evaluation metrics in the verification component of the RMS. The QMS should be adaptable to every type of AI system, ensuring that technical tests can be conducted on each AI system. The Strategy Pattern [6] (p. 506), a behavioral software design pattern, has been used to ensure a flexible design. Fig. 3 depicts the implementation design of the RMS verification component.
Figure 3: Architecture of the RMS Verification Component
The Strategy Pattern contains a context and a strategy. The strategy is abstract and contains several children with concrete strategy implementations. Then, the context can choose the best-fitting strategy for its task at runtime. The applied Strategy Pattern is modified with minor changes in the RMS verification component. The context is also an abstract class; concrete context child classes implement the context logic. The Model class is the abstract context class and represents the AI systems. Only the concrete context class LargeLanguageModel exists, but additional, such as a Linear Regression Model class, can be easily extended. The abstract Model class is connected to the abstract strategy classes: Performance, Explainability, and Consistency via a composition, meaning that the abstract strategy can only exist as long as the model class exists. The Model class contains a Performance, Explainability, and Consistency object in its constructor, which are initialized and further defined in its child class LargeLanguageModel. Specifically, the LargeLanguageModel class contains LLMPerformance, LLMExplainability, and LLMConsistency objects as concrete strategies in its constructor. The concrete strategy classes contain methods for each technical evaluation metric. Generally, this implementation design guarantees modularity and makes it easy to add new concrete contexts (types of AI systems), strategies (AI characteristics), and concrete strategies (technical evaluation metrics) without changing existing code.
3.4 Prototype QMS - First Version
The prototype QMS is accessible at https://power.bpm.cit.tum.de/qmsAIA/.
Main Service: Quality Management System. The user accesses the QMS home page after signing up and signing in. The prototype QMS home page design is depicted in Fig. 4.